A child's bank contains $6.30 in dimes and quarters. There are twice as many dimes as quarters. How many of each kind of coin are in the bank?
Answer by richwmiller(17219)

d=2q
10d+25q=630
10*2q+25q=630
20q+25q=630
45q=630
q=14
d=28
check
14*25=250+100=350
10*28=280
350+280=630
ok